OpenStudy (anonymous):

Write the complex number, z=-5+5i in polar form with the angle expressed in radians from 0 to 2(pi).

OpenStudy (dumbcow):

\[-5 = rcos \theta\] \[5 = rsin \theta\] solve for angle by substitution \[5 = -5\frac{\sin \theta}{\cos \theta}\] \[\tan \theta = -1\] \[\theta = \frac{3\pi}{4}\] \[r = \frac{5}{\sin(3\pi/4)} = \frac{5}{\sqrt{2}/2} = 5\sqrt{2}\]

OpenStudy (dumbcow):

there is an easier way as well using triangle  then use pythagorean thm to find r tangent to find angle

OpenStudy (anonymous):

sqrt of 50

OpenStudy (dumbcow):

which simplifies to 5sqrt2